# Grounded in Literature

### v1

@@ F.1 Additional Results and Significance Analysis
ScholarEval Baseline MeanSE (SD) MeanBL (SD) $p$ Sig.?
ScholarEval $_{\text{Llama-70B}}$ Llama-3.3-70B 2.04 (1.19) 1.81 (0.90) 0.094 No
ScholarEval $_{\text{Llama-70B}}$ GPT-4.1 2.04 (1.19) 2.14 (1.04) 0.460 No
ScholarEval $_{\text{Llama-70B}}$ Claude-4-Sonnet 2.04 (1.19) 2.01 (1.01) 0.827 No
ScholarEval $_{\text{Llama-70B}}$ o4-mini-deep-research 2.04 (1.19) 2.08 (1.06) 0.772 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.61 (1.42) 1.81 (0.90) 1.3e-06 Yes
ScholarEval $_{\text{GPT-4.1}}$ GPT-4.1 2.61 (1.42) 2.14 (1.04) 0.003 Yes
ScholarEval $_{\text{GPT-4.1}}$ Claude-4-Sonnet 2.61 (1.42) 2.01 (1.01) 1.2e-04 Yes
ScholarEval $_{\text{GPT-4.1}}$ o4-mini-deep-research 2.61 (1.42) 2.08 (1.06) 5.5e-04 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Llama-3.3-70B 2.64 (1.40) 1.81 (0.90) 3.7e-07 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ GPT-4.1 2.64 (1.40) 2.14 (1.04) 0.001 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Claude-4-Sonnet 2.64 (1.40) 2.01 (1.01) 5.2e-05 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ o4-mini-deep-research 2.64 (1.40) 2.08 (1.06) 2.4e-04 Yes
@@
ScholarEval Baseline MeanSE (SD) MeanBL (SD) $p$ Sig.?
ScholarEval $_{\text{Llama-70B}}$ Llama-3.3-70B 1.94 (1.13) 1.74 (0.87) 0.054 No
ScholarEval $_{\text{Llama-70B}}$ GPT-4.1 1.94 (1.13) 2.13 (1.00) 0.083 No
ScholarEval $_{\text{Llama-70B}}$ Claude-4-Sonnet 1.94 (1.13) 2.11 (1.03) 0.126 No
ScholarEval $_{\text{Llama-70B}}$ o4-mini-deep-research 1.94 (1.13) 2.35 (1.11) 3.8e-04 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.52 (1.51) 1.74 (0.87) 2.2e-09 Yes
ScholarEval $_{\text{GPT-4.1}}$ GPT-4.1 2.52 (1.51) 2.13 (1.00) 0.003 Yes
ScholarEval $_{\text{GPT-4.1}}$ Claude-4-Sonnet 2.52 (1.51) 2.11 (1.03) 0.002 Yes
ScholarEval $_{\text{GPT-4.1}}$ o4-mini-deep-research 2.52 (1.51) 2.35 (1.11) 0.212 No
ScholarEval $_{\text{Claude-4-Sonnet}}$ Llama-3.3-70B 2.90 (1.39) 1.74 (0.87) 8.1e-20 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ GPT-4.1 2.90 (1.39) 2.13 (1.00) 1.6e-09 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ Claude-4-Sonnet 2.90 (1.39) 2.11 (1.03) 9.3e-10 Yes
ScholarEval $_{\text{Claude-4-Sonnet}}$ o4-mini-deep-research 2.90 (1.39) 2.35 (1.11) 2.6e-05 Yes
@@
o4-mini-deep-research 2.98 (0.90) 2.12 (1.04) 2.17 (1.05) 2.60 (1.05) 2.41 (1.05) 2.08 (1.06)
@@
o4-mini-deep-research 3.33 0.49
@@ F.2 Manual Audit
These errors in attribution are also made by GPT-4o-search-preview. In the example below, it wrongly attributes a publication by Brotherton & Balskus (2013) to Kazane et al.
@@
Appendix G Additional Details from the Expert User Study
G.1 Setup and Materials
Rubric. We include our full rubric for the user study in Table 15. Answers were collected via Google Forms.
@@ G.1 Setup and Materials
G.2 Statistical Methods
Because we collect multiple research ideas per person, our data is no longer independent, and we cannot use a mean-differences t-test. Instead, we use a Linear Mixed-Effects Model (Raudenbush & Bryk, 2002), which models both fixed effects (ScholarEval vs deep research) and random effects (ideas within and across participants). This helps to stabilize the ratings across our experts, as some may be consistently higher or lower raters. This additionally helps to account for some variability from resarch idea quality. The Linear Mixed-Effects Model is defined as follows:
@@ G.2 Statistical Methods
Appendix H ScholarEval output examples
@@
Appendix I LLM Usage
The authors would like to declare the usage of LLMs for some aspects of code generation, LaTeX formatting and minor cosmetic improvements to the manuscript writing. However, LLMs were not used in the ideation of the project, designing the ScholarEval framework, or major content writing.
Marwa Abdulhai et al., “Defining Deception in Decision Making”
Philipp Guevorguian et al., “Exploring the Recall of Language Models: Case Study on Molecules”
Seungwon Oh et al., “Recovering Plasticity of Neural Networks via Soft Weight Rescaling”
Yuwei Yan et al., “OpenCity: A Scalable Platform to Simulate Urban Activities with Massive LLM Agents”
Moritz Glaser et al., “ESMGain: Effective and Efficient Prediction of Mutation’s functional Effect via ESM2 Transfer Learning and robust Benchmarks”
Svetlana Pavlova et al., “Flow Matching for One-Step Sampling”
Sasan Tavakkol et al., “Less is More: Adaptive Coverage for Synthetic Training Data”
Lorenzo Pacchiardi et al., “100 instances is all you need: predicting LLM success by testing on a few instances”
Zhihan Zhou et al., “GenomeOcean: Efficient Foundation Model for Genome Generation”
Eduardo Sánchez et al., “Linguini: A benchmark for language-agnostic linguistic reasoning”
Zekun Li et al., “MMSci: A Dataset for Graduate-Level Multi-Discipline Multimodal Scientific Understanding”
Alexander Shypula et al., “Does Instruction Tuning Reduce Diversity? A Case Study Using Code Generation”
Naman Gupta et al., “MAC-CAFE: Multi-actor, Centralized Critic Architecture for Feedback-driven Editing”
Magnus Müller et al., “Large-Scale Multi-Agent Reinforcement Learning for Traffic Signal Optimization”
Aohan Sun et al., “Mitigating Privacy Risk of Adversarial Examples with Counterfactual Explanations”
Gwok-Waa Wan et al., “GenBen: A Genarative Benchmark for LLM-Aided Design”
Kyeongrok Park et al., “Toward Human-Interpretable Explanations in a Unified Framework for GNNs”
Wenjie Tang et al., “StarCraft II Arena: Evaluating LLMs in Strategic Planning, Real-Time Decision Making, and Adaptability”
Chen Gao et al., “EmbodiedCity: A Benchmark Platform for Embodied Agent in Real-world City Environment”
Sungmin Han et al., “Improving Transformer Interpretability with Activation Contrast-Based Attribution”
Hong Xie et al., “Multiple-play Stochastic Bandits with Prioritized Resource Sharing”
Santiago Yeomans et al., “From Abstract Noise to Architectural Form: Designing Diffusion Models for Efficient Floor Plan Generation”
Chinmay Mittal et al., “FCoReBench: Can Large Language Models Solve Challenging First-Order Combinatorial Reasoning Problems?”
Haihong Yang et al., “scKGOT: Intercellular Signaling Inference with Knowledge Graph Optimal Transport for Single-cell Transcriptomics”
Changliang Zhou et al., “ICAM: Rethinking Instance-Conditioned Adaptation in Neural Vehicle Routing Solver”
Yisheng Xiao et al., “Path Selection Makes BERT-family Good Generators”
Xiang Liu et al., “ChunkKV: Semantic-Preserving KV Cache Compression for Efficient Long-Context LLM Inference”
Teng Yan et al., “Vision-Based Pseudo-Tactile Information Extraction and Localization for Dexterous Grasping”
Pavel Strashnov et al., “Towards Robust Evaluation of Protein Generative Models: A Systematic Analysis of Metrics”
Zhenlei Wang et al., “Robust Heterogeneous Treatment Effect Estimation under Covariate Perturbation”
Yao Shiyi et al., “BID: Broad Incremental for Android Malware Detection”
Leo McKee-Reid et al., “Honesty to Subterfuge: In-Context Reinforcement Learning Can Make Honest Models Reward Hack”
Tingzhou Wei et al., “Multivariate Time-series Forecasting with SPACE: Series Prediction Augmented by Causality Estimation”
Anuradha Kumari et al., “ZEPHYR GAN: REDEFINING GAN WITH FLEXIBLE GRADIENT CONTROL”
Andrei Chertkov et al., “Tensor Train Decomposition for Adversarial Attacks on Computer Vision Models”
Zhenghan Chen et al., “Advancing Drug-Target Interaction Prediction via Graph Transformers and Residual Protein Embeddings”
Zhenghan Chen et al., “Non-Commutative Spectral Geometry for Adaptive Quantum-Classical Drug-Target Interaction Prediction”
Chris Cameron et al., “Foundation Models for Boolean Logic”
Haoxuan Li et al., “Principle Counterfactual Fairness”
Yuntian Wu et al., “Invariant Spatiotemporal Representation Learning for Cross-patient Seizure Classification”
Benas et al., “Modeled grid cells aligned by a flexible attractor”
Wittkamp et al., “The neural dynamics of positive and negative expectations of pain”
Cui et al., “Dysfunctional S1P/S1PR1 signaling in the dentate gyrus drives vulnerability of chronic pain-related memory impairment”
O’Leary et al., “Natural forgetting reversibly modulates engram expression in hippocampal feedforward circuits”
Klaassen et al., “Basolateral amygdala inhibition impairs updating of appetitive and aversive values by interacting with the prefrontal cortex”
Haupt et al., “The transformation of sensory to perceptual braille letter representations in the visually deprived brain”
Liu et al., “Cell class-specific long-range axonal projections of neurons in mouse whisker-related somatosensory cortices”
Campbell et al., “Human single-neuron activity is modulated by intracranial theta burst stimulation of the basolateral amygdala”
Derkaloustian et al., “Fine Touch Perception Relies on Frictional Instabilities”
Lee et al., “The influence of temporal context on vision over multiple time scales”
Setogawa et al., “Acquisition of auditory discrimination mediated by different processes through two distinct circuits linked to the lateral striatum”
Cooper et al., “Ultraslow serotonin oscillations in the hippocampus delineate substates across NREM and waking”
Mollá–Albaladejo et al., “Molecular characterization of gustatory second-order neurons reveals integrative mechanisms of gustatory and metabolic information”
Bloem et al., “Dynamic estimation of the attentional field from visual cortical activity”
Xu et al., “Neural Representation of Time across Complementary Reference Frames”
Rieser et al., “Multifaceted Role of Galanin in Whole Brain Excitability”
Lu et al., “The interplay between homeostatic synaptic scaling and homeostatic structural plasticity maintains the robust firing rate of neural networks”
Ecker et al., “Assemblies, synapse clustering and network topology interact with plasticity to explain structure–function relationships of the cortical connectome”
Dash et al., “Rules for reactivation across REM sleep microstates following sensory fear learning”
March et al., “The Hungry Lens: Hunger Shifts Attention and Attribute Weighting in Dietary Choice”
Molkov et al., “Introducing perturbations in point-process models of excitable systems”
Huang et al., “Neural coding of multiple motion speeds in visual cortical area MT”
Liu et al., “Striatal Crosstalk Between Dopamine and Serotonin Systems”
Kang et al., “Rapid rebalancing of co-tuned ensemble activity in the auditory cortex”
Praegel et al., “Age and Learning Shapes Sound Representations in Auditory Cortex During Adolescence”
Zhang et al., “Oxytocin restores context-specific hyperaltruistic preference”
Hall et al., “A cortical–hippocampal communication undergoes rebalancing after new learning”
Zhang et al., “Humans underestimate their body mass in microgravity”
Barnby et al., “Self–other generalisation shapes social interaction and is disrupted in borderline personality disorder”
Tardiff et al., “Normative evidence weighing and accumulation in correlated environments”
Wu et al., “The Self-Interest of Adolescents Overrules Cooperation in Social Dilemmas”
Chen et al., “Synchronous Ensembles of Hippocampal CA1 Pyramidal Neurons During Novel Exploration”
Wang et al., “The relationship between cognitive abilities and mental health as represented by cognitive abilities at the neural and genetic levels of analysis”
Zhong et al., “Modular DNA Barcoding of Nanobodies Enables Multiplexed in situ Protein Imaging and High-throughput Biomolecule Detection”
Majhi et al., “Non-autonomous cell redox-pairs dictate niche homeostasis in multi-lineage stem populations”
Krwawicz et al., “Introduction of cytosine-5 DNA methylation sensitizes cells to oxidative damage”
Xiu et al., “Action mechanism of a novel agrichemical quinofumelin against Fusarium graminearum
Chang et al., “Cancer cells differentially modulate mitochondrial respiration to alter redox state and enable biomass synthesis in nutrient-limited environments”
Mohanty et al., “Deep Learning Reveals Endogenous Sterols as Allosteric Modulators of GPCRs”
Wang et al., “Structure and evolution of Alanine/Serine Decarboxylases & S-Adenosylmethionine Decarboxylases in plants”
Wei et al., “Crystal structure and catalytic mechanism of PL35 family glycosaminoglycan lyases with an ultrabroad substrate spectrum”
Govorunova et al., “Blue-shifted ancyromonad channelrhodopsins for multiplex optogenetics”
He et al., “Coordinated regulation of chemotaxis and resistance to copper by CsoR”
Jandu et al., “Membrane mimetic thermal proteome profiling ($M$-TPP) enables proteome-wide target engagement in membranes”
Chong et al., “Establishing the foundations for a data-centric AI approach for virtual drug screening”
Schulze et al., “Effects of residue substitutions on the cellular abundance of proteins ”
Maus et al., “Screening the MMV Pathogen Box reveals the mitochondrial bc1-complex as a drug target in mature Toxoplasma gondii bradyzoites”
Lefroncois et al., “The Role of ATP Synthase Subunit e (ATP5I) in Mediating the Metabolic and Antiproliferative Effects of Biguanides”
[…]
Marks et al., “Determining the off-target activity of antibiotics and novel translation initiation sites in mitochondria”
Leanza et al., “Increased bone inflammation in type 2 diabetes and obesity correlates with Wnt signaling downregulation and reduced bone strength”
Zhang et al., “Distinct mechanisms of inhibition of Kv2 potassium channels by tetraethylammonium and RY785”
Luo et al., “Isobaric crosslinking mass spectrometry technology for studying conformational and structural changes in proteins and complexes”
Antenucci et al., “Reassessing the substrate specificities of the major Staphylococcus aureus peptidoglycan hydrolases lysostaphin and LytM”
Zhou et al., “Structural insights into human propionyl-CoA carboxylase …”
Liu et al., “Genome-wide mapping of native co-localized G4s and R-loops …”
D’Oliveira et al., “Recognition and Cleavage of Human tRNA …”
Rucci et al., “Effects of blood meal source and seasonality on reproductive traits of Culex quinquefasciatus (Diptera: Culicidae)”
García-Ruiz et al., “Fitness drivers of division of labor in vertebrates”
Nakagawa et al., “An illusion of a macroecological law, abundance–occupancy relationship in birds”
Jiang et al., “Assessing plant phenological changes based on drivers of spring phenology”
Howard–Spink et al., “Old age variably impacts chimpanzee engagement and efficiency in stone tool use”
Rebindaine et al., “Developmental constraints mediate the summer solstice reversal of climate effects on European beech bud set”
Smit et al., “Risk-taking incentives predict aggression heuristics in female gorillas”
Croijmans et al., “Strip cropping shows promising increases in ground beetle community diversity compared to monocultures”
Wang et al., “Loss of olfaction reduces caterpillar performance and increases susceptibility to a natural enemy”
Tao et al., “Partitioning changes in ecosystem productivity by effects of species interactions in biodiversity experiments”
Yang et al., “Interpreting prediction intervals and distributions for biologically meaningful effects”
Gao et al., “Pesticide-induced resurgence in brown planthoppers is mediated by action on a suite of genes that promote juvenile hormone biosynthesis …”
Fargeot et al., “Genetic diversity affects ecosystem functions across trophic levels …”
Seltzer et al., “Female Moths Incorporate Plant Acoustic Emissions into Their Oviposition Decision-Making Process”
Seguchi et al., “Vasopressin 1a receptor antagonist disrupts male–male affiliative relationships formed by triadic cohabitation in large-billed crows”
Gatt et al., “Integrating microscopy and transcriptomics from individual eukaryotic plankton (Ukiyo-e-Seq)”
Rydhmer et al., “Automating an insect biodiversity metric using distributed optical sensors: an evaluation across Kansas, USA cropping systems”
Diaz–Colunga et al., “Full factorial construction of synthetic microbial communities”
Zhang et al., “Neuropeptide bursicon and its receptor mediate the transition in seasonal polyphenism of Cacopsylla chinensis
Zhang et al., “Birds migrate longitudinally in response to the resultant Asian monsoons of the Qinghai–Tibet Plateau uplift”

### v2

@@ G.1 Additional Results and Significance Analysis
[…]
ScholarEval $_{\text{Llama}}$ Llama-3.3-70B 1.94 (1.13) 1.74 (0.87) 0.027 Yes
GPT-4.1 1.94 (1.13) 2.13 (1.00) 0.958 No
GPT-5.1 Instant 1.94 (1.13) 1.95 (0.97) 0.537 No
Claude-4-Sonnet 1.94 (1.13) 2.11 (1.03) 0.937 No
GPT-4o-search-preview 1.94 (1.13) 1.95 (1.02) 0.536 No
OpenAI Deep Research 1.94 (1.13) 2.35 (1.11) 1.000 No
DR Tulu 1.94 (1.13) 2.39 (1.20) 1.000 No
ScholarEval $_{\text{GPT-4.1}}$ Llama-3.3-70B 2.52 (1.51) 1.74 (0.87) <<0.001 Yes
GPT-4.1 2.52 (1.51) 2.13 (1.00) 0.002 Yes
GPT-5.1 Instant 2.52 (1.51) 1.95 (0.97) <<0.001 Yes
Claude-4-Sonnet 2.52 (1.51) 2.11 (1.03) 0.001 Yes
GPT-4o-search-preview 2.52 (1.51) 1.95 (1.02) <<0.001 Yes
OpenAI Deep Research 2.52 (1.51) 2.35 (1.11) 0.106 No
DR Tulu 2.52 (1.51) 2.39 (1.20) 0.177 No
ScholarEval $_{\text{GPT-5.1}}$ Llama-3.3-70B 2.09 (1.28) 1.74 (0.87) <<0.001 Yes
GPT-4.1 2.09 (1.28) 2.13 (1.00) 0.633 No
GPT-5.1 Instant 2.09 (1.28) 1.95 (0.97) 0.115 No
Claude-4-Sonnet 2.09 (1.28) 2.11 (1.03) 0.567 No
GPT-4o-search-preview 2.09 (1.28) 1.95 (1.02) 0.120 No
OpenAI Deep Research 2.09 (1.28) 2.35 (1.11) 0.982 No
DR Tulu 2.09 (1.28) 2.39 (1.20) 0.991 No
ScholarEval $_{\text{Claude}}$ Llama-3.3-70B 2.90 (1.39) 1.74 (0.87) <<0.001 Yes
GPT-4.1 2.90 (1.39) 2.13 (1.00) <<0.001 Yes
GPT-5.1 Instant 2.90 (1.39) 1.95 (0.97) <<0.001 Yes
Claude-4-Sonnet 2.90 (1.39) 2.11 (1.03) <<0.001 Yes
GPT-4o-search-preview 2.90 (1.39) 1.95 (1.02) <<0.001 Yes
OpenAI Deep Research 2.90 (1.39) 2.35 (1.11) <<0.001 Yes
DR Tulu 2.90 (1.39) 2.39 (1.20) <<0.001 Yes
@@
OpenAI Deep Research 2.98 (0.90) 2.12 (1.04) 2.17 (1.05) 2.60 (1.05) 2.41 (1.05) 2.08 (1.06)
@@
System Coverage (mean $\pm$ std)
OpenAI Deep Research $2.35\pm 1.07$
ScholarEval $_{\text{GPT}}$ $2.84\pm 1.39$
ScholarEval $_{\text{Claude}}$ $2.91\pm 1.34$
ScholarEval $_{\text{o4-mini}}$ $2.78\pm 1.35$
Table 15: ScholarEval coverage on ScholarIdeas-AI using o4-mini as backbone
System Coverage (mean (std))
Llama-3.3-70B 2.22 (1.04)
GPT-4.1 2.78 (1.15)
Claude-4-Sonnet 2.57 (1.05)
GPT-4o-search-preview 2.36 (1.05)
OpenAI Deep Research 2.66 (1.06)
ScholarEval-Llama-3.3 2.52 (1.17)
ScholarEval-GPT-4.1 3.45 (1.34)
ScholarEval-Claude-4-Sonnet 3.34 (1.34)
Idea Novelty Checker (Shahid et al., 2025) 2.27 (1.15)
Table 16: Coverage of contribution rubrics in ScholarIdeas-AI of Idea Novelty Checker (Shahid et al., 2025)
@@
OpenAI Deep Research 3.33 0.49
@@ G.2 Manual Audit
These errors in attribution are also made by GPT-4o-search-preview. In the example below, it wrongly attributes a publication by Brotherton and Balskus (2013) to Kazane et al.
@@
Appendix H Additional Details from the Expert User Study
H.1 Setup and Materials
Rubric We include our full rubric for the user study in Table 18. Answers were collected via Google Forms.
@@ H.1 Setup and Materials
H.2 Statistical Methods
Because we collect multiple research ideas per person, our data is no longer independent, and we cannot use a mean-differences t-test. Instead, we use a Linear Mixed-Effects Model (Raudenbush and Bryk, 2002), which models both fixed effects (ScholarEval vs deep research) and random effects (ideas within and across participants). This helps to stabilize the ratings across our experts, as some may be consistently higher or lower raters. This additionally helps to account for some variability from resarch idea quality. The Linear Mixed-Effects Model is defined as follows:
@@ H.2 Statistical Methods
Appendix I ScholarEval output examples
@@
List of AI papers in ScholarIdeas.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
  20. 20.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
  25. 25.
  26. 26.
  27. 27.
  28. 28.
  29. 29.
  30. 30.
  31. 31.
  32. 32.
  33. 33.
  34. 34.
  35. 35.
  36. 36.
  37. 37.
  38. 38.
  39. 39.
  40. 40.
List of neuroscience papers in ScholarIdeas.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
  20. 20.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
  25. 25.
  26. 26.
  27. 27.
  28. 28.
  29. 29.
  30. 30.
  31. 31.
  32. 32.
  33. 33.
List of biochemistry papers in ScholarIdeas.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
  20. 20.
  21. 21.
  22. 22.
  23. 23.
  24. 24.
List of ecology papers in ScholarIdeas.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
  13. 13.
  14. 14.
  15. 15.
  16. 16.
  17. 17.
  18. 18.
  19. 19.
  20. 20.